### v5

@@ 4 Refusal option
This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects. Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is employed.
@@
Suppose that the modified uniform rank-minimizing mechanism is employed. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominates $r_{a}$ with (or without) the refusal option.
By adopting $f^{M}$, any ODS can fail because increasing the number of acceptable objects may lead the agent to be unassigned to truly acceptable objects, depending on the strategy profile of the other agents. Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS. However, $f^{M}$ may lead to other strategic problems. That is, if $f^{M}$ is adopted instead of $f^{U}$, an agent may have an incentive to reduce the number of acceptable object types.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$. Let $\delta=\left(o_{1},o_{2},\varnothing\right)$ and $\delta^{\prime}=\left(o_{1},\varnothing,o_{2}\right)$. We assume $q_{o}=1$. Suppose $r_{a_{1}}=\delta$ and $p_{a_{2}}=\delta$. In this case, since $\left(\delta(=r_{a_{1}}),\delta\right)$ does not satisfy all of (i), (ii) and (iii),
$\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2},$
$\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
@@
$\left(f^{M}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ 1, ($1,\text{ }\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=0,$
$\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
Thus, in this case, if $f^{M}$ is employed, then $a_{1}$ has an incentive to choose $\delta^{\prime}$ instead of $r_{a_{1}}=\delta$. In other words, in this case, $a_{1}$ has an incentive to reveal by reducing its truly acceptable types. On the other hand, if $f^{U}$ is employed, then $a_{1}$ has no such an incentive.
Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a_{1}}$. However, agents may have such incentives even if they do not have the refusal option.
@@ 5 Concluding Remarks
We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then truth-telling is strategically dominated by an ODS. The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
@@
Therefore, if resolving the problem in that way is difficult, it is necessary to consider alternative mechanisms. First, as shown in this study, if we modify the uniform rank-minimizing mechanism, then truth-telling is not strategically dominated by any other strategies, including the outside option demotion ones. However, in this modification, it is necessary to ensure that increasing the number of acceptable types can result in a disadvantage. Thus, this modification may lead to the opposite problem, where agents strategically decrease the number of acceptable types.
@@
$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}x_{a\bar{o}}$ $<$ $1$
@@
$Y=\left\{y\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{ao}=1\right.\right\},Y^{\prime}=\left\{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{a^{\prime}o}^{\ast}=1\right.\right\}.$
@@
We consider an arbitrary integer $k\geq 3$. Then, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$. We assume that under $\left(r_{a},p_{-a}\right)$ the agents are divided into $k$ groups. For the first group, there are $q_{r_{a}^{1}}+1$ agents whose preference is equivalent to $r_{a}$ under $\left(r_{a},p_{-a}\right)$; that is, if $a^{\prime}$ is in this group, then $r_{a}=p_{a^{\prime}}$. Note that $a$ belongs to this group. Next, we consider $j$th group for each $j=2,3,\cdots,k-1$. In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose first-best object type is $r_{a}^{j}$ and the second-best object is $r_{a}^{j+1},$ the third-best $r_{a}^{j+2},$cdots and the $k$th-best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$. If $a_{j}$ belongs $j$th group for $j=1,2,\cdots,k-1$, the preference orders of them are as follows:
@@
Let $f$ be a fair mechanism, $p_{a}\neq r_{a}$ and $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. First, suppose $k\leq\bar{k}\left(p_{a}\right)$. By Lemma 2, there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$. Therefore, in this case, $r_{a}$ is not strategically dominate by $p_{a}$.
Second, suppose $k>\bar{k}\left(p_{a}\right)$. In this case, since $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$. Therefore, $p_{a}$ does not strategically dominate $r_{a}$.
@@
We use the example introduced in the proof of Theorem 2. Suppose that there are $a\in\mathcal{A}$, $o,o^{\prime}\in\mathcal{O}$ satisfying (3). Let $d_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$ and $p_{a^{\prime}}=d_{a}$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Then, since (3) is satisfied and $f^{U}$ satisfies ETE, there is $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, $y_{ao^{\prime}}=1$ and $y_{a^{\prime}o^{\prime\prime}}=1$ for some $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $o^{\prime\prime}\in\mathcal{O}$ such that $R\left(o^{\prime},d_{a}\right)<R\left(o^{\prime\prime},d_{a}\right)$. Then, if $y$ is realized, $a$ refuses $o^{\prime}$ but $a^{\prime}$ prefers $o^{\prime}$ to $o^{\prime\prime}$. Therefore, $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
@@
First, if $r_{a}^{1}=\varnothing,$ then $\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{a\varnothing}=1$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Thus, in this case, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$. Thus, we assume $r_{a}^{1}\neq\varnothing$. We consider $p_{a}\left(\neq r_{a}\right)$. Let $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$.
@@
Therefore, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$.

### v6

@@ 4 Refusal option
This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects. Therefore, $r_{a_{1}}$ is not dominated by $d_{a_{1}}$, if $f^{M}$ is employed.
@@
For $f^{M}$ and $a$ with any $r_{a}\in\mathcal{P}$, truth-telling is not a dominated strategy with the refusal option.
By adopting $f^{M}$, any ODS can fail because increasing the number of acceptable objects may lead the agent to be unassigned to truly acceptable objects, depending on the strategy profile of the other agents. Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS.
However, the modified URM may introduce alternative strategic concerns. We now turn to a different class of strategic manipulations.
Let $\Pi_{a}\subseteq\mathcal{P}$ be a set of preferences such that $\pi_{a}\in\Pi_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,\pi_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,\pi_{a}\right)>R\left(\varnothing,\pi_{a}\right)$, (ii) $R\left(\varnothing,\pi_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (iii) $R\left(o,r_{a}\right)=R\left(o,\pi_{a}\right)+1$ for all $o$ such that $R\left(\varnothing,\pi_{a}\right)\leq R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (iv) $R\left(o,r_{a}\right)=R\left(o,\pi_{a}\right)$ for all $o$ such that $R\left(\varnothing,r_{a}\right)<R\left(o,r_{a}\right)$. We call the strategy of agent $a$ of revealing a preference in $\Pi_{a}$ an outside option promotion strategy (hereafter OPS) of agent $a$. In words, contrary to ODSs, an agent raises the rank of the outside option in their reported preference, making fewer object types appear acceptable.
First, we have the following result on this class of strategies on the URM (not modified one).
Proposition 6
For $f^{M}$ and $a$ with any $r_{a}\in\mathcal{P}$, any OPS of $a$ is weakly dominated by truth-telling of $a$. Moreover, this is satisfied even without the refusal option.
On the other hand, if the modified URM is adopted instead of the URM, an agent may have an incentive to adopt an OPS. We briefly show this fact by using the following example.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$. Let $\delta=\left(o_{1},o_{2},\varnothing\right)$, $\delta^{\prime}=\left(o_{1},\varnothing,o_{2}\right)$ and $\delta^{\prime\prime}=\left(\varnothing,o_{1},o_{2}\right)$. We assume $q_{o_{1}}=q_{o_{2}}=1$. Suppose $r_{a_{1}}=\delta$, $p_{a_{2}}=\delta$ and $p_{a_{3}}=\delta^{\prime\prime}$. Then, $\delta^{\prime}$ is an OPS of $a_{1}$. In this case, since $\left(\delta(=r_{a_{1}}),\delta\right)$ does not satisfy all of (i), (ii) and (iii),
$\left(f^{M}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{M}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=\frac{1}{2},$
$\left(f^{U}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
@@
$\left(f^{M}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $1,\text{ }\left(f^{M}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=0,$
$\left(f^{U}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
Thus, in this case, if $f^{M}$ is employed, then $a_{1}$ has an incentive to choose $\delta^{\prime}$ instead of $r_{a_{1}}=\delta$. In other words, in this case, $a_{1}$ has an incentive to use an OPS. On the other hand, if $f^{U}$ is employed, then $a_{1}$ has no such an incentive.
Formally, we have the following result.
Proposition 7
Let $f$ be a fair rank-minimizing rule such that, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that dominates $r_{a}$ with a refusal option. Then, there is some $(r_{a},p_{-a})$ such that $a$ strictly prefers $g\left(f\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f\left(r_{a},p_{-a}\right),r\right)$ where $\pi_{a}$ is an OPS of $a$. Moreover, this is satisfied even without the refusal option; that is, for that $(r_{a},p_{-a})$, $a$ strictly prefers $f\left(\pi_{a},p_{-a}\right)$ to $f\left(r_{a},p_{-a}\right)$.
Proposition 5 shows that it is possible to construct a fair rank-minimizing mechanism under which truth-telling is not a dominated strategy. However, doing so gives rise to a new type of strategic concern that does not exist under the URM. Specifically, as shown in Proposition 6, under the URM, any OPS is dominated by truth-telling, thereby eliminating the risk of such manipulation. In contrast, Proposition 7 implies that under a fair rank-minimizing mechanism in which truth-telling is not dominated by any strategies, including the modified URM, an OPS is not dominated by truth-telling; that is, for an agent, there exsits some profile of other agents This suggests that manipulation through such strategies may persist even in mechanisms designed to eliminate strategic dominance of truth-telling.
Of course, this does not mean that truth-telling is dominated by an OPS. However, this reveals a trade-off inherent in fair rank-minimizing mechanisms: while modifying them can eliminate the strategic dominance of truth-telling, such modifications may introduce incentives for other manipulations.
@@ 5 Concluding Remarks
We show that if an assignment is determined by the URM, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then truth-telling is dominated by an ODS. The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
@@
Therefore, if resolving the problem in that way is difficult, it is necessary to consider alternative mechanisms. First, as shown in this study, if we modify the URM, then truth-telling is not dominated by any other strategies, including the outside option demotion ones. However, in this modification, it is necessary to ensure that increasing the number of acceptable types can result in a disadvantage. Thus, this modification leads to the opposite strategic problem, where agents have an incentive to decrease the number of acceptable types.
@@
$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}x_{a\bar{o}}$ $<$ $1.$
@@
[…]
@@
We consider an arbitrary integer $k\geq 3$. Then, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$. We assume that under $\left(r_{a},p_{-a}\right)$ the agents are divided into $k$ groups. For the first group, there are $q_{r_{a}^{1}}+1$ agents whose preference is equivalent to $r_{a}$ under $\left(r_{a},p_{-a}\right)$; that is, if $a^{\prime}$ is in this group, then $r_{a}=p_{a^{\prime}}$. Note that $a$ belongs to this group. Next, we consider $j$th group for each $j=2,3,\cdots,k-1$. In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose first-best object type is $r_{a}^{j}$ and the second-best object is $r_{a}^{j+1},$ the third-best $r_{a}^{j+2},$ $\cdots$ and the $k$th-best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$. If $a_{j}$ belongs $j$th group for $j=1,2,\cdots,k-1$, the preference orders of them are as follows:
@@
Let $f$ be a fair mechanism, $p_{a}\neq r_{a}$ and $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. First, suppose $k\leq\bar{k}\left(p_{a}\right)$. By Lemma 2, there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$. Therefore, in this case, $r_{a}$ is not dominate by $p_{a}$.
Second, suppose $k>\bar{k}\left(p_{a}\right)$. In this case, since $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$. Therefore, $p_{a}$ does not dominate $r_{a}$.
@@
Let $r_{a}=\left(o_{1},\varnothing,\cdots\right)$ and $r_{a^{\prime}}=\left(o_{1},o_{2},\varnothing,\cdots\right)$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Furthermore, let $d=\left(d_{\bar{a}}\right)_{\bar{a}\in\mathcal{A}}$ such that $d_{\bar{a}}=(o_{1},o_{2},o_{3},\cdots,\varnothing)$ for all $\bar{a}\in\mathcal{A}$; that is, $d$ is a profile in which every agent adopts an ODS.
First, we show that no agent has an incentive to unilaterally deviate from $d$. It is sufficient to show that for all $\bar{a}\in\mathcal{A}$ and all $p_{\bar{a}}\in\mathcal{P}$, either
$\left(f^{U}\left(d\right)\right)_{\bar{a}o_{1}}>\left(f^{U}\left(p_{\bar{a}},d_{{}_{-\bar{a}}}\right)\right)_{\bar{a}o_{1}}$ (18)
or
$\left(f^{U}\left(d\right)\right)_{\bar{a}\bar{o}}=\left(f^{U}\left(p_{\bar{a}},d_{{}_{-\bar{a}}}\right)\right)_{\bar{a}\bar{o}},\,\ \text{for all }\bar{o}\in\mathcal{O},$ (19)
because $o$ is the best object type for all agents.
Let $k$ be the smallest integer such that $p_{\bar{a}k}\neq d_{\bar{a}k}$. By Lemma 2, if $k>\bar{k}\left(p_{\bar{a}}\right)$, then (19) is satisfied. We assume $k\leq\bar{k}\left(p_{{}_{\bar{a}}}\right)$; that is,
$\sum\limits_{\bar{o}:R(\bar{o},p_{\bar{a}})<k}q_{\bar{o}}\left(=\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$
is satisfied for that $k$. Let $o^{\prime\prime}$ such that $R\left(o^{\prime\prime},p_{\bar{a}}\right)=k<R\left(o^{\prime\prime},d_{\bar{a}}\right)$. Let $y\in\mathcal{Y}^{\ast}\left(d\right)$ such that $y_{\bar{a}o_{k}}=1$. Moreover, we arbitrarily choose $y^{\prime}$ such that $y_{\bar{a}o_{1}}^{\prime}=1$.
We consider two cases. First, we assume
$\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})\leq k}q_{\bar{o}}=\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})<k}q_{\bar{o}}+q_{o_{k}}<\left|\mathcal{A}\right|.$
In this case, some agent is assigned to $o_{k+1}$ in a rank-minimizing assignment under $d$. Let $y^{\prime\prime}$ be such that $y_{\bar{a}o^{\prime\prime}}^{\prime\prime}=1$, $y_{a^{\prime}o_{k}}^{\prime\prime}=1$ for $a^{\prime}$ with $y_{a^{\prime}o_{k+1}}=1,$ and $y_{a^{\prime\prime}\bar{o}}^{\prime\prime}=y_{a^{\prime\prime}\bar{o}}$ for all $a^{\prime\prime}\in\mathcal{A\setminus}\left\{\bar{a},a^{\prime}\right\}$ and $\bar{o}\in\mathcal{O}$. Then,
$RV(y^{\prime\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right))<RV(y,d)\leq RV(y^{\prime},d)\leq RV(y^{\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right)).$
Thus, $y^{\prime}\notin\mathcal{Y}^{\ast}\left(p_{\bar{a}},d_{-\bar{a}}\right)$ and (18) is satisfied in this case.
Second, we assume
$\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})\leq k}q_{\bar{o}}=\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})<k}q_{\bar{o}}+q_{o_{k}}\geq\left|\mathcal{A}\right|.$
Then, since $q_{o_{1}}<\left|\mathcal{A}\right|$, $k>1$. Let $y^{\prime\prime}$ be such that $y_{\bar{a}o^{\prime\prime}}^{\prime\prime}=1$, and $y_{a^{\prime\prime}\bar{o}}^{\prime\prime}=y_{a^{\prime\prime}\bar{o}}$ for all $a^{\prime\prime}\in\mathcal{A\setminus}\left\{\bar{a}\right\}$ and $\bar{o}\in\mathcal{O}$. Moreover, let $y^{\prime\prime\prime}$ be an arbitrary assignment in $\mathcal{Y}^{\ast}\left(d\right)$ such that $y_{\bar{a}o_{k^{\prime}}}^{\prime\prime\prime}=1$ for some $k^{\prime}=1,\cdots,k-1$.
$RV(y^{\prime\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right))=RV(y,d)=RV(y^{\prime\prime\prime},d)=RV(y^{\prime\prime\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right)).$
Therefore, (19) is satisfied in this case.
Finally, suppose $q_{o_{1}}+q_{o_{2}}<\left|\mathcal{A}\right|$. We show that $g\left(f^{U}\left(d\right),r\right)$ is wasteful. Since $q_{o_{1}}+q_{o_{2}}<\left|\mathcal{A}\right|$, there is $y\in\mathcal{Y}^{\ast}\left(d\right)$, $y_{ao_{2}}=1$ and $y_{a^{\prime}o_{1}}=y_{a^{\prime}o_{2}}=0$ for some $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Thus, $\left(f^{U}\left(d\right)\right)_{ao_{2}}>0$ and $\left(g\left(f^{U}\left(d\right),r\right)\right)_{ao_{2}}=0$. Therefore, $g\left(f^{U}\left(d\right),r\right)$ is wasteful.
@@
First, if $r_{a}^{1}=\varnothing,$ then $\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{a\varnothing}=1$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Thus, in this case, there is no $p_{a}\in\mathcal{P}$ that dominate $r_{a}$. Thus, we assume $r_{a}^{1}\neq\varnothing$. We consider $p_{a}\left(\neq r_{a}\right)$. Let $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$.
@@
Therefore, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that dominates $r_{a}$.
Proof of Proposition 6
Suppose not; that is, there exist $(r_{a},p_{-a})\in\mathcal{P}^{\left|\mathcal{A}\right|}$ such that $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$ where $\pi_{a}$ is an OPS of $a$. By the definition of OPS, when $o$ and $o^{\prime}$ satisfy $R\left(o,r_{a}\right)<R\left(o^{\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and $R\left(o^{\prime},\pi_{a}\right)>R\left(\varnothing,\pi_{a}\right),$
$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $=$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\text{ if }y_{ao}=1,$ (20)
$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $<$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\text{ if }y_{a\varnothing}=1,$ (21)
$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $>$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\text{ if }y_{ao^{\prime}}=1.$ (22)
First, we show that $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=0$. Suppose not; that is, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ for some $o$ satisfying $R\left(\varnothing,r_{a}\right)\leq R\left(o,r_{a}\right)$. By Lemma 2, $y_{a\varnothing}=1$. Then, by (20), (21) and (22), for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $y_{a\varnothing}^{\prime}=1$. However, this contradicts that $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. Therefore, $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=0$; that is, for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\varnothing}=0$.
Then, for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}=1$ implies $R\left(\varnothing,r_{a}\right)>R\left(o,r_{a}\right)$. Since $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$, for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $y_{a\varnothing}^{\prime}=0$ is also satisfied.
Next, we show that for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}=1$ implies $R\left(o,\pi_{a}\right)<R\left(\varnothing,\pi_{a}\right)$. Suppose not; that is, $y_{ao}=1$ implies $R\left(o,\pi_{a}\right)>R\left(\varnothing,\pi_{a}\right)$. Then, let $y^{\prime}$ be such that $y_{a\varnothing}^{\prime}=1$ and $y_{\bar{a}\bar{o}}^{\prime}=y_{\bar{a}\bar{o}}$ for any other $\left(\bar{a},\bar{o}\right).$ By (20), (21) and (22),
$RV\left(y^{\prime},\left(\pi_{a},p_{-a}\right)\right)$ $<$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\leq$
$RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$ $\leq$ $RV\left(y^{\prime\prime},\left(\pi_{a},p_{-a}\right)\right),$
for any $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=0$, because $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. This contradicts that for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $y_{a\varnothing}^{\prime}=0$. Thus, for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}=1$ implies $R\left(o,\pi_{a}\right)<R\left(\varnothing,\pi_{a}\right)$.
We arbitrarily choose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. By (20),
$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $=$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\leq$
$RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$ $\leq$ $RV\left(y^{\prime\prime},\left(\pi_{a},p_{-a}\right)\right),$
for any $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=0$. Since $y_{a\varnothing}^{\prime}=0$ for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)=\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$ but this contradicts that $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
The proof above is satisfied even without the refusal option, because $g\left(f\left(\cdot,\cdot\right),r\right)=f\left(\cdot,\cdot\right)$ is satisfied for all $g\left(f\left(\cdot,\cdot\right),r\right)$ above.
Proof of Proposition 7
Let $f$ be a fair rank-minimizing rule such that, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that dominates $r_{a}$ with the refusal option. First, we consider the following specific example. Let $a_{1},a_{2},a_{3}\in\mathcal{A}$ and $o_{1},o_{2}\in\mathcal{O}$. Suppose $q_{o_{1}}=q_{o_{2}}=1$. Let $\delta=\left(o_{1},\varnothing,\cdots\right)$ and $\delta^{\prime}=\left(o_{1},o_{2},\cdots,\varnothing\right)$. First, we assume $r_{a_{1}}=p_{a_{2}}=\delta$, $R\left(o_{2},p_{a_{3}}\right)=1,$ and $R\left(\varnothing,p_{a}\right)$ for any $a\in\mathcal{A\setminus}\left\{a_{1},a_{2},a_{3}\right\}$ if there exists some. Then, $\delta^{\prime}$ is an ODS of $a,$ and
$\left(f\left(\delta,p_{-a_{1}}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\delta,p_{-a_{1}}\right)\right)_{a_{1}\varnothing}=\frac{1}{2}\text{,}$
$\left(f\left(\delta^{\prime},p_{-a_{1}}\right)\right)_{a_{1}o_{1}}$ $=$ $1$
Therefore, $a_{1}$ strictly prefers $g\left(f\left(\delta^{\prime},p_{-a_{1}}\right),r\right)$ to $g\left(f\left(\delta,p_{-a_{1}}\right),r\right)$.
Since $\delta^{\prime}$ does not dominate $r_{a_{1}}$ under $f$, there is some $p_{-a_{1}}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a_{1}$ strictly prefers $g\left(f\left(\delta,p_{-a_{1}}^{\prime}\right),r\right)$ to $g\left(f\left(\delta^{\prime},p_{-a}^{\prime}\right),r\right)$. Next, we assume that $r_{a_{1}}=\delta^{\prime}$. Then, $\delta$ is an OPS of $a_{1}$. Therefore, $a_{1}$ strictly prefers $g\left(f\left(\pi_{a},p_{-a_{1}}^{\prime}\right),r\right)$ to $g\left(f\left(r_{a_{1}},p_{-a}^{\prime}\right),r\right)$, where $r_{a_{1}}=\delta^{\prime}$ and $\pi_{a}=\delta$.
The proof above is satisfied even without the refusal option, because $g\left(f\left(\cdot,\cdot\right),r\right)=f\left(\cdot,\cdot\right)$ is satisfied for all $g\left(f\left(\cdot,\cdot\right),r\right)$ above.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author used ChatGPT (OpenAI) in order to improve readability and language. After using this, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.